O O 11. Among the four methods discussed in this chapter, the table method was considered the most desirable. 12. Other thingTRUE/FALSE, Need #s11-20
Answer #1

1.
False, financial calculator/excel/formula are more desirable

2.
True

3.
True
FV=100*1.1^5=161.0510000

4.
True

5.
False
PV=500/1.11^8=216.9632481

6.
True

7.
True

8.
False

9.
False
r=(200/100)^(1/20)-1=3.53%

10.
True
n=log(200/100)/log(1+7%)=10.2447684